# Review of Demographic Bias in Face Recognition

### v1

@@ II-D Image Quality
The quality of input images and associated covariates significantly influence the manifestation of demographic bias in FR systems. Numerous studies have emphasized how disparities in image quality across different demographic groups can lead to variations in system performance. For instance, Cavazos et al. [24] analyzed both data-driven and scenario-based factors, revealing that dataset complexity and decision thresholds have a notable impact on recognition accuracy and racial bias. Their experiments across multiple algorithms further demonstrated that East Asian faces required higher decision thresholds compared to Caucasian faces to achieve equivalent error rates, highlighting the interplay between dataset characteristics and demographic attributes.
@@
The study conducted by MdTF highlighted skin reflectance as a critical factor influencing both the accuracy and efficiency of FR systems [38, 39]. Analyzing 158 FR systems, they found that lower skin reflectance, typically associated with darker skin-tones, correlated with reduced accuracy and higher transaction times. These effects varied across systems, underscoring the role of image acquisition quality as a stronger predictor of performance as mentioned earlier. Similarly, Wu et al. [40] explored the effects of brightness and illumination, demonstrating that under-exposed or over-exposed images result in higher FMRs, while significant brightness differences between image pairs diminish similarity scores. They recommended controlled image acquisition processes to achieve consistent brightness across demographic groups, thereby reducing accuracy disparities.
@@
Krishnapriya et al. [41] further examined how variations in image quality contributed to performance gaps between African-American and Caucasian cohorts. Enhancing image quality notably reduced these disparities, particularly by minimizing low-similarity errors within the genuine distribution. Following ICAO compliance guidelines, they evaluated biometric sample quality to support these findings. Albiero et al. [36] investigated gender-based disparities in FR systems, linking these to differences in genuine and imposter score distributions. They also identified confounding factors such as cosmetics and image pose. Despite using neutral and balanced datasets, their study revealed that such measures alone were insufficient to fully eliminate observed disparities.
@@
Work by Lu et al. provided a detailed analysis of the influence of covariates on FR performance, incorporating variables such as skin tone, age, gender, pose, facial hair, and occlusion across three datasets and five FR systems [32]. Their findings highlighted that skin tone significantly affects verification accuracy, with lighter skin tones consistently outperforming medium-dark tones. However, they also emphasized the challenges posed by ambiguities in skin tone classification, advocating for more precise methodologies for performance assessments. In alignment with earlier studies, Lu et al. observed that male subjects generally achieved better recognition accuracy than female subjects. They attributed this disparity to factors such as occlusion caused by longer hair and alterations in facial appearance due to makeup. These observations corroborate prior findings indicating that facial makeup can negatively impact recognition accuracy [46, 47]. Collectively, these studies underscore that demographic bias in FR systems is intrinsically tied to image quality and related covariates, necessitating focused efforts to address these issues systematically.
@@ II-E Combined or Intersectional Factors
In the preceding sections, we examined the individual factors contributing to demographic bias in FR. This section shifts focus to studies that investigate the combined effects of multiple demographic attributes, such as age, race, and gender. Vera-Rodriguez et al. [42] emphasized the significance of gender as a covariate in FR, observing that males consistently outperform females across various demographic groups. These findings highlight the necessity of addressing combined demographic factors to achieve equitable outcomes.
@@
Age-related biases have been linked to structural transformations in facial features over time, particularly among children. Ricanek et al. [37] observed increased complexity of child aging compared to adults, attributing recognition challenges to changes in facial bone structure and the proportions of facial components. Additionally, Best-Rowden and Jain [43] reported nuanced patterns in age-related recognition performance, noting that while males generally exhibit higher genuine scores, their performance declines more rapidly with age compared to females. These observations underline the intricate interplay of demographic attributes in shaping biases in FR systems. The intersection of age, race, and gender significantly amplifies biases in FR systems. In [44], Sarridis et al. identified a disproportionately high mistreatment rate for African females over 60 years compared to Caucasians, illustrating the compounded effects of intersecting demographic factors. Similarly, El Khiyari et al. [45] demonstrated that face verification accuracy is notably lower for younger individuals (aged 18–30), females, and certain racial groups such as Black individuals, highlighting the challenges posed by such intersections of demographic factors.
@@
Algorithmic evaluations further reinforce these findings. The FRVT report [15] observed elevated false positives among children and the elderly, particularly within Asian and American Indian groups. These disparities were intensified in low-quality imaging conditions, with younger and older demographics experiencing higher error rates. Cook et al. [39] extended this understanding by showing that self-reported demographic factors like age and measured skin lightness significantly impact recognition scores, often compounded by environmental factors such as illumination. Collectively, these studies underscore the intricate challenges of addressing intersecting biases in FR systems, emphasizing the necessity of age-specific and multi-faceted considerations in algorithm design and evaluation.
III Datasets for Demographic Bias
@@
In evaluating and mitigating demographic bias in FR systems, the selection of suitable datasets plays an important role. Although numerous FR datasets exist, those specifically intended for bias and fairness-related tasks must include demographic labels associated with each subject or identity. The datasets designed for tasks such as race, gender, ethnicity, or age estimation are particularly useful when they include demographic labels, as such factors are critical for assessing fairness. For certain tasks (related to estimation or classification of attributes), having a single image per subject may suffice, as training and testing can be conducted separately for each image. However, FR models, especially state-of-the-art systems, benefit from having multiple images (variations) of each identity to train more robust feature extractors. The testing phase (verification or identification) requires multiple images per identity, where one image serves as the gallery/ template and the others are used as test or probe samples. Occasionally, the gallery is also composed of more than one image per subject.
@@
Dataset Name Year Number of Images / Subjects Demographic Labels Typical Purpose
@@
TABLE II: Commonly used datasets for tasks realted to demographic bias in face recognition.
These requirements significantly reduce the availability of datasets suitable for assessing demographic bias, as most FR datasets do not provide adequate demographic labels or have highly skewed distributions of subjects across different demographic groups. In this section, we outline recent and commonly used publicly available datasets that are relevant for assessing and addressing demographic bias in FR systems.
  • MORPH [48]: The MORPH is one of the largest facial image datasets available in several variants and versions. MORPH-II is the most commonly referred academic version, comprising more than 55,000 images from more than 13,000 subjects. Despite its usefulness, it should be noted that the dataset is highly skewed in terms of gender and ethnicity, with a significant over-representation of male subjects (more than 46,000 images) and a limited number of female subjects (approximately 8,500 images). Race labels are also provided, with categories including Black, White, Asian, Hispanic, and others. However, due to the demographic imbalance, this dataset may introduce bias in FR tasks focused on fairness and equity.
@@
  • VGGFace2 [51]: The VGGFace2 is a large-scale FR dataset containing over 3.31 million images of 9,131 subjects. This dataset was annotated for gender (Male, Female) and ethnicity (Asian, Black, Indian, White) labels by Idiap Research Institute [^3] Annotations for VGGFace2 Dataset making it useful for bias related tasks. The dataset has been noted to have a bias towards White and male subjects, which should be considered when using it for fairness and bias studies. The access to original download location has been removed by its creators as of 2024.
  • DemogPairs [52]: DemogPairs is a validation set containing 10.8K images, divided into six demographic folds: Asian females, Asian males, Black females, Black males, White females, and White males. The dataset was specifically designed to evaluate the demographic bias in FR models, offering 58.3 million evaluation pairs, including cross-demographic, cross-gender, and cross-ethnicity pairs. The DemogPairs dataset was constructed with rigorous demographic annotation and is a useful resource for testing the generalization of FR systems across diverse demographic groups.
  • Racial Faces in-the-Wild (RFW) [27]: RFW is a benchmarking dataset designed to study racial bias in FR systems. It consists of four subsets (African, Asian, Caucasian, and Indian), each containing about 3,000 individuals and 6,000 image pairs (these pairs have been defined by its creators). The dataset is specifically used for face verification tasks and includes balanced pairs of genuine (mated) and imposter (non-mated) images. The RFW dataset has been widely adopted by the research community to evaluate and compare the performance of FR algorithms across different racial groups.
  • BUPT-BalancedFace [53]: The BUPT-BalancedFace dataset was constructed to address demographic bias by ensuring race balance across the dataset. It contains approximately 1.3 million images from 28,000 celebrities, with a balanced distribution of 7,000 identities per race. The dataset was selected from MS-Celeb-1M [60] through the FreeBase and Face++ APIs, although it has been noted that the labels may contain noise. Due to its size and balanced nature, the BUPT-BalancedFace dataset has become a popular resource for training and fine-tuning FR models while mitigating race-related biases.
@@
  • DiveFace [54]: DiveFace is a dataset generated from the Megaface dataset (now decommissioned) [61], containing over 120,000 images from 24,000 identities. The dataset includes two gender and three ethnicity classes, allowing for detailed demographic analysis. Annotations were made using a semi-automatic process, followed by manual inspection. This dataset is useful for studying the impact of gender and ethnicity in FR tasks, although it may exhibit bias in some groups. It should also be noted that in DiveFace dataset, the subjects of Indian and African ethnicities have been grouped together– which can make it difficult to use it in conjunction with other datasets that typically do not follow such grouping.
@@
  • Multiple Encounter Dataset (MEDS-II) [55]: MEDS-II is an extension of the MEDS-I dataset and was created to assist with the NIST Multiple Biometric Evaluation. The dataset includes over 1,300 images of 518 subjects, with many subjects having only a single image, limiting its usefulness for verification tasks. The MEDS-II is dominated by male subjects of White and Black ethnicities. Despite its limitations in demographic diversity, it remains a useful resource for testing FR systems in real-world scenarios, especially where multiple encounters of a subject are available.
@@
  • BFW (Balanced Faces in the Wild) [56] [57]: The BFW dataset was designed to provide a more balanced evaluation of FR systems by creating subgroups that are evenly split across gender and ethnicity. The dataset is compiled from VGGFace2 [51] and offers a refined approach to subgroup analysis with less overlap between training and testing data. The corresponding demographic labels were generated using ethnicity [62] and gender [63] classifiers, followed by manual validation. Additionally, the BFW dataset is also balanced with respect to the number of images, subjects, and the (ratio of) images per subject; making it particularly useful for evaluating the demographic fairness of FR models, offering a more balanced alternative to other datasets.
@@
  • CASIA-Face-Africa [58]: The CASIA-Face-Africa dataset is the first large-scale face dataset of African subjects– comprising 38,546 images from 1,183 individuals, captured under varying illumination conditions using multi-spectral cameras. It includes detailed demographic attributes and facial expressions along with manually annotated with facial key points. The dataset exhibits a well-distributed age representation, with a significant portion belonging to the subjects upto 40 years, aligning with the majority workforce demographics. Additionally, it maintains an almost balanced gender ratio (48% male, 52% female), making it useful for geneder-based analysis as well. In terms of ethnic variations, the dataset includes multiple African ethnic groups, with a notable dominance of the Hausa ethnic group.
@@
  • CausalFace [59]: It is a large-scale dataset of synthetically generated faces comprising 48,000 synthetic face image pairs generated from 10,200 unique identities, along with 555,000 human annotations covering individual attributes and pairwise identity comparisons. The dataset is constructed using EG3D [64], a state-of-the-art GAN framework that allows explicit control over geometry and pose. In terms of demographic attributes, the CausalFace dataset includes six demographic groups created by combining three ethnicities (White, Black, and East Asian) and two genders (male and female). Using a GAN-based generator, face images were systematically modified across unprotected attributes such as pose, age, expression, and lighting while maintaining identity consistency. The dataset images were selected on the basis of perceived matching scores provided by human annotators.
@@
The datasets are summarized in Table II, while Fig. 2 illustrates an overview of the demographic distribution, with race and gender as demographic variables, for the datasets reviewed in this section.
IV Assessment of Demographic Bias in FR

### v2

@@ II-D Image Quality
The quality of input images and associated covariates significantly influence the manifestation of demographic differentials in FR systems. Numerous studies have emphasized how disparities in image quality across different demographic groups can lead to variations in system performance. For instance, Cavazos et al. [24] analyzed both data-driven and scenario-based factors, revealing that dataset complexity and decision thresholds have a notable impact on recognition accuracy and racial bias. Their experiments across multiple algorithms further demonstrated that East Asian faces required higher decision thresholds compared to Caucasian faces to achieve equivalent error rates, highlighting the interplay between dataset characteristics and demographic attributes.
@@
The study conducted by MdTF highlighted skin reflectance as a critical factor influencing both the accuracy and efficiency of FR systems [40, 41]. Analyzing 158 FR systems, they found that lower skin reflectance, typically associated with darker skin-tones, correlated with reduced accuracy and higher transaction times. These effects varied across systems, underscoring the role of image acquisition quality as a stronger predictor of performance as mentioned earlier. Similarly, Wu et al. [42] explored the effects of brightness and illumination, demonstrating that under-exposed or over-exposed images result in higher FMRs, while significant brightness differences between image pairs diminish similarity scores. They recommended controlled image acquisition processes to achieve consistent brightness across demographic groups, thereby reducing accuracy disparities. In [43], authors analyzed the effect of facial brightness on recognition performance disparities between Caucasian and African American females. Their experiments showed that by balancing brightness–using either median pixel values or full pixel distributions–the similarity scores improve for both groups.
@@
Krishnapriya et al. [44] further examined how variations in image quality contributed to performance gaps between African-American and Caucasian cohorts. Enhancing image quality notably reduced these disparities, particularly by minimizing low-similarity errors within the genuine distribution. Following ICAO compliance guidelines, they evaluated biometric sample quality to support these findings. Albiero et al. [38] investigated gender-based disparities in FR systems, linking these to differences in genuine and imposter score distributions. They also identified confounding factors such as cosmetics and image pose. Despite using neutral and balanced datasets, their study revealed that such measures alone were insufficient to fully eliminate observed disparities. Bhatta et al.  [45] explored 1-to-many identification performance under degraded image conditions resembling surveillance-quality probes. The findings showed that blur and low resolution significantly increased false matches, with demographic disparities being more pronounced between genders than racial groups, underscoring the influence of image quality on recognition fairness. The investigation of demographic fairness in the context of 1-to-many (1:$N$) face identification has been comparatively under-explored. Another recent study in area [59] observed that the accuracy variation patterns in 1:$N$ identification are similar to those reported in 1:1 verification scenarios. Furthermore, their analysis highlighted that degradations in image quality–particularly due to blur or reduced resolution–tend to increase the false positive identification rate (FPIR).
@@
Work by Lu et al. provided a detailed analysis of the influence of covariates on FR performance, incorporating variables such as skin-tone, age, gender, pose, facial hair, and occlusion across three datasets and five FR systems [34]. Their findings highlighted that skin tone significantly affects verification accuracy, with lighter skin tones consistently outperforming medium-dark tones. However, they also emphasized the challenges posed by ambiguities in skin-tone classification, advocating for more precise methodologies for performance assessments. In alignment with earlier studies, Lu et al. observed that male subjects generally achieved better recognition accuracy than female subjects. They attributed this disparity to factors such as occlusion caused by longer hair and alterations in facial appearance due to makeup. These observations corroborate prior findings indicating that facial makeup can negatively impact recognition accuracy [60, 61]. Collectively, these studies underscore that demographic fairness in FR systems is intrinsically tied to image quality and related covariates, necessitating focused efforts to address these issues systematically.
@@ II-E Combined or Intersectional Factors
In the preceding sections, we examined the individual factors contributing to demographic differentials in FR. This section shifts focus to studies that investigate the combined effects of multiple demographic attributes, such as age, race, and gender. Vera-Rodriguez et al. [46] emphasized the significance of gender as a covariate in FR, observing that males consistently outperform females across various demographic groups. These findings highlight the necessity of addressing combined demographic factors to achieve equitable outcomes.
@@
Age-related biases have been linked to structural transformations in facial features over time, particularly among children. Ricanek et al. [39] observed increased complexity of child aging compared to adults, attributing recognition challenges to changes in facial bone structure and the proportions of facial components. Additionally, Best-Rowden and Jain [47] reported nuanced patterns in age-related recognition performance, noting that while males generally exhibit higher genuine scores, their performance declines more rapidly with age compared to females. These observations underline the intricate interplay of demographic attributes in shaping biases in FR systems. The intersection of age, race, and gender significantly amplifies accuracy differentials in FR systems. In [48], Sarridis et al. identified a disproportionately high number of incorrect verifications (both false positives and false negative)for African females over 60 years compared to Caucasians, illustrating the compounded effects of intersecting demographic factors. Similarly, El Khiyari et al. [49] demonstrated that face verification accuracy is notably lower for younger individuals (aged 18–30), females, and certain racial groups such as Black individuals, highlighting the challenges posed by such intersections of demographic factors. Liu et al.  [62] observed that multiple factors such as ethnicity, pose, occlusion, and blur jointly impact FR accuracy.
Among the few studies addressing explanatory frameworks, [50] introduced a methodology aimed at evaluating the influence of both protected and non-protected attributes on FR performance. Their analysis revealed that fairness outcomes exhibit considerable variability across multi-attribute groupings. Notably, patterns observed in single-attribute evaluations frequently diminish or invert when extended to analyses involving multiple attributes.
@@
Algorithmic evaluations further reinforce these findings. The FRVT report [15] observed elevated false positives among children and the elderly, particularly within Asian and American Indian groups. These disparities were intensified in low-quality imaging conditions, with younger and older demographics experiencing higher error rates. Cook et al. [41] extended this understanding by showing that self-reported demographic factors like age and measured skin lightness significantly impact recognition scores, often compounded by environmental factors such as illumination. However, more recent analyses, such as Albiero et al. [51], suggest that recognition accuracy may not consistently improve or decline with age. Their findings indicate that face matchers do not exhibit a simple age-performance trend, challenging assumptions from earlier works.
The inconsistent findings indicate that age-related bias in FR remains insufficiently understood and lacks a clear consensus. Overall these studies highlight the complexity of mitigating intersecting biases within FR systems, underscoring the importance of incorporating age-specific and multi-factor considerations in both algorithm development and evaluation protocols.
II-F Soft Attributes or Non-Demographic Factors
While prior sections have explored to demographic factors influencing disparities in FR performance, recent research has highlighted the critical role of non-demographic factors– such as facial hair, hairstyle, and makeup– in shaping recognition outcomes. These factors, though not inherently demographic, are deeply intertwined with social and cultural norms that vary across gender and ethnicity. In the context of FR, these soft attributes function as partial occlusions or lead to shifts in the underlying data distribution, thereby introducing additional challenges. It is well-documented that occlusions can significantly degrade recognition accuracy in biometric and generic vision applications [63, 64, 65, 66]. However, when these occlusions vary systematically across groups, they contribute to unequal recognition outcomes and further exacerbate existing disparities, mimicking demographic bias.
Recent studies have demonstrated that many of the observed demographic fairness in FR may in fact be driven by these correlated non-demographic traits. Through extensive experiments, Kurz et al. showed that the gender gap in recognition accuracy vanishes when males and females share the same facial attributes, indicating that social appearance conventions, rather than gender itself, are a primary factor [58]. Pangelinan et al. similarly found that controlling for facial pixel information across genders improved non-matching errors for females, pointing to the role of hairstyles and other aesthetic choices in accuracy differentials [54]. Multiple works have rigorously examined the impact of facial hair, showing that variations in facial hair significantly affect both impostor and genuine score distributions, and that these effects persist even with balanced training datasets[55, 56]. In females, factors such as hairstyle, facial morphology, and overall facial dimensions result in a reduced number of effective face pixels, which has been associated with elevated false match rates (FMR) compared to males [67]. Bhatta et al.  [53] demonstrated that balancing for hairstyle-induced occlusions in test data can nearly eliminate the gender gap. Additionally, Albiero et al. illustrated that recognition performance can vary across different facial regions, further suggesting that visual composition and occlusions, rather than demographic identity alone, play a pivotal role toward recognition [52]. Complementing these findings, Mamede et al. [57] proposed the Face Occlusion Impact Ratio (FOIR)- a novel metric to measure the influence of occlusions on recognition performance across demographic groups. Their results show that occlusions impact demographic groups unequally, with African individuals disproportionately affected- thereby amplifying existing biases.
These collective insights underscore the significant influence of non-demographic but demographically correlated appearance factors in shaping recognition performance. While these factors are not inherently biological, but rather linked to the demographic groups through social and cultural appearance norms and styles. Understanding and accounting for these nuanced sources of variation is essential to mitigate differential accuracy and build fair FR systems.
III Datasets for the Study of Demographic Accuracy Differences
@@
In evaluating and mitigating demographic fairness in FR systems, the selection of suitable datasets plays an important role. Although numerous FR datasets exist, those specifically intended for bias and fairness-related tasks must include demographic labels associated with each subject or identity. The datasets designed for tasks such as race, gender, ethnicity, or age estimation are particularly useful when they include demographic labels, as such factors are critical for assessing fairness. For certain tasks (related to estimation or classification of attributes), having a single image per subject may suffice, as training and testing can be conducted separately for each image. However, FR models, especially state-of-the-art systems, benefit from having multiple images (variations) of each identity to train more robust feature extractors. The testing phase or deployment has two primary modes: one-to-one (1:1) verification and one-to-many (1:$N$) identification. In 1:1 matching, a probe image is compared against one or more enrolled images of a claimed identity. For identification (1:$N$ matching), the probe is compared against a gallery containing reference images from multiple identities to determine the closest match. Occasionally, the gallery may be composed of multiple images per identity.
TABLE III: Commonly used datasets for tasks related to demographic accuracy differentials in face recognition. Appendix VII provides the URLs of the referenced datasets.
@@
Dataset Name Year # Images / Subjects Demographic Labels Typical Purpose
@@
MORPH-III [67] 2021 127,000+ / 26,401 (Male, Female), (Black, White) Train/ Test
@@
These requirements significantly reduce the availability of datasets suitable for assessing demographic fairness, as most FR datasets do not provide adequate demographic labels or have highly skewed distributions of subjects across different demographic groups. It is important to note that several prominent bias evaluation efforts, such as those led by NIST and the MdTF, rely on proprietary or restricted-access datasets. While these organizations have conducted large-scale, operationally relevant evaluations, their limited availability to the wider research community poses challenges for reproducibility, benchmarking, and comparative analysis. In this section, we outline recent and commonly used publicly available datasets that are relevant for assessing and addressing demographic fairness in FR systems.
  • MORPH [68]: The MORPH dataset is one of the largest facial image datasets available in several variants and versions. MORPH-II is the most commonly referred academic version, comprising more than 55,000 images from more than 13,000 subjects. Despite its usefulness, it should be noted that the dataset is highly skewed in terms of gender and ethnicity, with a significant over-representation of male subjects (more than 46,000 images) and a limited number of female subjects (approximately 8,500 images).
    The MORPH-III dataset has seen increased adoption in recent research investigating demographic variations in FR [67]. For improved data quality, the dataset was curated to remove duplicates, twins, and mislabeled entries, resulting in a refined set comprising 127,000+ images of nearly 26,000 subjects. This curated subset provides a cleaner and demographically structured resource for evaluating bias. For both versions, the datasets consist of mugshot-style facial images acquired under controlled conditions, featuring neutral expressions, approximately frontal poses, and consistent illumination. However, it has been noted to have demographic imbalances, particularly under-representation of certain groups (e.g., Asian, Indian), which may affect fairness studies.
@@
  • VGGFace2 [71]: The VGGFace2 dataset is a large-scale FR dataset containing over 3.31 million images of 9,131 subjects. This dataset was annotated for gender (Male, Female) and ethnicity (Asian, Black, Indian, White) labels by Idiap Research Institute [^1] Annotations for VGGFace2 Dataset making it useful for fairness related tasks. This dataset offers significant advantages due to a large number of identities with variations in age, pose, and lighting; VGGFace2 improves over the original with more diversity too. The dataset, however, is significantly imbalanced, with a disproportionate representation of White and male subjects. Additionally, the original download source has been removed by the creators as of 2024.
  • DemogPairs [72]: The DemogPairs dataset is a validation set containing 10.8K images, divided into six demographic folds: Asian females, Asian males, Black females, Black males, White females, and White males. The dataset was specifically designed to evaluate the demographic bias in FR models, offering 58.3 million evaluation pairs, including cross-demographic, cross-gender, and cross-ethnicity pairs. The DemogPairs dataset was constructed with rigorous demographic annotation and is a useful resource for testing the generalization of FR systems across diverse demographic groups.
  • Racial Faces in-the-Wild (RFW) [29]: The RFW dataset is a benchmarking dataset designed to study racial fairness in FR systems. It consists of four subsets (African, Asian, Caucasian, and Indian), each containing about 3,000 individuals and 6,000 image pairs (these pairs have been defined by its creators). The dataset is specifically used for face verification tasks and includes balanced pairs of genuine (mated) and imposter (non-mated) images. The RFW dataset has been widely adopted by the research community to evaluate and compare the performance of FR algorithms across different racial groups. However, given the overlap between the RFW dataset and the widely used training dataset, MS-Celeb-1M, it is important to consider that the evaluation set may have been seen by the model during training. The creators of RFW have also provided a list of non-overlapping images to address this issue.
  • BUPT-BalancedFace [73]: The BUPT-BalancedFace dataset was constructed to address demographic fairness by ensuring race balance across the dataset. It contains approximately 1.3 million images from 28,000 celebrities, with a balanced distribution of 7,000 identities per race. The dataset was selected from MS-Celeb-1M dataset [80] through the FreeBase and Face++ APIs, although it has been noted that the labels may contain noise. Due to its size and balanced nature, the BUPT-BalancedFace dataset has become a popular resource for training and fine-tuning FR models while mitigating race-related biases. Two key points should be considered: first, the demographic labels, which are based on nationality or API, contain some noise. Second, the dataset overlaps with MS-Celeb-1M dataset, which restricts certain experiments when using models trained exclusively on MS-Celeb-1M dataset.
@@
  • DiveFace [74]: The DiveFace is a dataset generated from the Megaface dataset (now decommissioned) [81], containing over 120,000 images from 24,000 identities. The dataset includes two gender and three ethnicity classes, allowing for detailed demographic analysis. Annotations were made using a semi-automatic process, followed by manual inspection. This dataset is useful for studying the impact of gender and ethnicity in FR tasks, although it may exhibit bias in some groups. It should also be noted that in DiveFace dataset, the subjects of Indian and African ethnicities have been grouped together– which can make it difficult to use it in conjunction with other datasets that typically do not follow such grouping.
@@
  • Multiple Encounter Dataset (MEDS-II) [75]: The MEDS-II dataset is an extension of the MEDS-I dataset and was created to assist with the NIST Multiple Biometric Evaluation. The dataset includes over 1,300 images of 518 subjects, with many subjects having only a single image, limiting its usefulness for verification tasks. The MEDS-II dataset is dominated by male subjects of White and Black ethnicities. Despite its limitations in demographic diversity, it remains a useful resource for testing FR systems in real-world scenarios, especially where multiple encounters of a subject are available. Additionally, the dataset is considerably smaller compared to current benchmarks.
@@
  • BFW (Balanced Faces in the Wild) [76] [77]: The BFW dataset was designed to provide a more balanced evaluation of FR systems by creating subgroups that are evenly split across gender and ethnicity. The dataset is compiled from VGGFace2 [71] and offers a refined approach to subgroup analysis with less overlap between training and testing data. The corresponding demographic labels were generated using ethnicity [82] and gender [83] classifiers, followed by manual validation. Additionally, the BFW dataset is also balanced with respect to the number of images, subjects, and the (ratio of) images per subject; making it particularly useful for evaluating the demographic fairness of FR models, offering a more balanced alternative to other datasets. However, this dataset shares the same two limitations as some earlier ones: a smaller size and overlap with VGGFace2.
@@
  • CASIA-Face-Africa [78]: The CASIA-Face-Africa dataset is the first large-scale face dataset of African subjects– comprising 38,546 images from 1,183 individuals, captured under varying illumination conditions using multi-spectral cameras. It includes detailed demographic attributes and facial expressions along with manually annotated with facial key points. The dataset exhibits a well-distributed age representation, with a significant portion belonging to the subjects up to 40 years, aligning with the majority workforce demographics. Additionally, it maintains an almost balanced gender ratio (48% male, 52% female), making it useful for gender-based analysis as well. In terms of ethnic variations, the dataset includes multiple African ethnic groups, with a notable dominance of the Hausa ethnic group.
@@
  • CausalFace [79]: It is a large-scale dataset of synthetically generated faces comprising 48,000 synthetic face image pairs generated from 10,200 unique identities, along with 555,000 human annotations covering individual attributes and pairwise identity comparisons. The dataset is constructed using EG3D [84], a state-of-the-art GAN framework that allows explicit control over geometry and pose. In terms of demographic attributes, the CausalFace dataset includes six demographic groups created by combining three ethnicities (White, Black, and East Asian) and two genders (male and female). Using a GAN-based generator, face images were systematically modified across unprotected attributes such as pose, age, expression, and lighting while maintaining identity consistency. The dataset images were selected on the basis of perceived matching scores provided by human annotators. Although it offers a bias-free ground truth, its synthetic nature may limit generalizability to real-world scenarios.
@@
The datasets are summarized in Table III, and Fig. 2 illustrates an overview of the demographic distribution, with race and gender as demographic variables, for the datasets reviewed in this section.
IV Assessment of Demographic Fairness in FR